When 18.0 mL of a 1.43×10-4 M silver nitrate solution is combined with 22.0 mL of a 4.94×10-4 M potassium hydroxide solution does a precipitate form? (yes or no) For these conditions the Reaction Quotient, Q, is equal to=
Answer #1

Answer

No

Q = 1.75 ×10-8

Explanation

Concentration of Ag+ after mixing the solution = 1.43×10-4M/(40ml/18ml) = 6.44×10-5M

Concentration of OH- after mixing the solution = (4.94 ×10-4M/(40/22ml) = 2.72×10-4M

Q = [Ag+][OH-]

Q = 6.44× 10-5M × 2.72 ×10-4

Q = 1.75×10-8

Ksp of AgOH = 2.0 ×10-8

Q < Ksp

Therefore

precipitate will not form